OpenStudy (anonymous):

Find the LCM of the set of polynomials. 121x^2-4y^2, 11x^2+2yx

OpenStudy (anonymous):

\[121x^{2}-4y ^{2}, 11x ^{2}+2yx\]

OpenStudy (anonymous):

first one is the difference of two squares. it factors as \[121x^2-4y^2=(11x+2y)(11x-2y)\] second factors as \[11x^2+2xy=x(11x+2y)\]

OpenStudy (anonymous):

you are looking for the LCM so take each factor you see and multiply them together. you only need the \(11x+2y\) once, so take \[x(11x+2y)(11x-2y)\] as LCM

OpenStudy (anonymous):

oh okay, thanks so much!

OpenStudy (anonymous):

yw